OpenStudy (anonymous):

Helpppppppppppppp Let f(x)=4x^3+7x^2-13x-3 and g(x)=x+3. Find f(x)/g(x)

OpenStudy (diyadiya):

\[\frac{f(x)}{g(x)} = \frac{4x^3+7x^2-13x-3}{x+3}\]

jhonyy9 (jhonyy9):

4x3 +7x2 -13x -3 / x+3 =4x2 -5x +2 4x3 +12x2 ------------ 0 -5x2 -13x -5x2 -15x ------------ 0 2x -3 2x +6 ------- 0 -9

jhonyy9 (jhonyy9):

hope that is understandably

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

A Lil Can you re write it clearly

jhonyy9 (jhonyy9):

f(x)=g(x)*(4x2 -5x+2) -9

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

whats the solution ? f(x)/g(x) is ?

jhonyy9 (jhonyy9):

is 4x2 -5x +2 and remainder -9